# Iterative Distillation for Reward-Guided Fine-Tuning of Diffusion Models in Biomolecular Design

### v2

@@ Abstract
We address the problem of fine-tuning diffusion models for reward-guided generation in biomolecular design. While diffusion models have proven highly effective in modeling complex, high-dimensional data distributions, real-world applications often demand more than high-fidelity generation, requiring optimization with respect to potentially non-differentiable reward functions such as physics-based simulation or rewards based on scientific knowledge. Although RL methods have been explored to fine-tune diffusion models for such objectives, they often suffer from instability, low sample efficiency, and mode collapse due to their on-policy nature. In this work, we propose an iterative distillation-based fine-tuning framework that enables diffusion models to optimize for arbitrary reward functions. Our method casts the problem as policy distillation: it collects off-policy data during the roll-in phase, simulates reward-based soft-optimal policies during roll-out, and updates the model by minimizing the KL divergence between the simulated soft-optimal policy and the current model policy. Our off-policy formulation, combined with KL divergence minimization, enhances training stability and sample efficiency compared to existing RL-based methods. Empirical results demonstrate the effectiveness and superior reward optimization of our approach across diverse tasks in protein, small molecule, and regulatory DNA design. The source code is released at (https://github.com/divelab/VIDD).
@@ 1 Introduction
To address the aforementioned challenges, we propose a new framework, VIDD (Value-guided Iterative Distillation for Diffusion models), designed to maximize possibly non-differentiable rewards in a stable and effective manner. The core idea is to iteratively distill soft-optimal policies—including, which serve as target denoising processes that optimize the reward while remaining close to the current fine-tuned model, as visualized in Figure 1. Concretely, the algorithm proceeds in three iterative steps: (1) roll-in using sufficiently exploratory off-policy trajectories, (2) roll-out to simulate soft-optimal policies, and (3) update the fine-tuned model by minimizing the KL divergence between the soft-optimal and current model policies. Importantly, in (2) and (3), our algorithm effectively leverages value functions tailored to diffusion models to guide fine-tuning, analogous to value-weighted MLE in RL (Peters et al., 2010). Notably, in contrast to methods such as PPO, our framework leverages off-policy roll-ins—decoupling data collection from policy updates—and employs forward KL minimization rather than reverse KL, both of which contribute to improved optimization stability over complex reward landscapes.
Our contributions are summarized as follows. We propose a novel algorithm, VIDD, for fine-tuning diffusion models through iterative distillation of target policies composed of both value functions and the current policy. Unlike direct reward backpropagation methods, which require differentiable reward signals, our approach can handle non-differentiable rewards commonly encountered in scientific domains. Compared to PPO-based methods, our algorithm achieves more stable optimization. Algorithmically, our key innovation lies in effectively incorporating value functions specifically tailored to the diffusion models. Empirically, we validate the effectiveness of our method across a range of scientific tasks, particularly in protein and small molecular design.
@@ 1.1 Related Works
When reward functions are differentiable, as is often the case in computer vision, state-of-the-art methods achieve strong performance by directly backpropagating gradients induced by the reward functions on the diffusion model (Clark et al., 2023; Prabhudesai et al., 2023; Wu et al., 2024; Wang et al., 2024). However, in many scientific domains such as biology and molecule field, reward functions are often non-differentiable (Lisanza et al., 2024; Hie et al., 2022). In such settings, a natural approach is to adopt RL techniques that optimize reward without requiring differentiability (Zhang et al., 2024; Gupta et al., 2025; Ektefaie et al., 2024; Vogt et al., 2024; Deng et al., 2024; Venkatraman et al., 2024; Rector-Brooks et al., 2024; Zekri and Boullé, 2025). For instance, DPOK (Fan et al., 2023) and DDPO (Black et al., 2024) adapt PPO, a stabilized variant of policy gradient methods, to the diffusion model fine-tuning. However, these methods often suffer from instability and slow convergence when applied to diffusion models (Clark et al., 2023, Figure 3), due to their inherently online nature and the use of a reverse KL divergence objective between the fine-tuned policy and a target policy (discussed further in Section 5). In contrast, our method avoided on-policy updates and instead leverages a forward KL divergence objective, which stabilizes training and mitigates mode collapse.
@@
An alternative line of work focuses on training-free techniques that aim to improve rewards solely at inference time. The most straightforward example is Best-of-N sampling, which selects the highest-reward output from N generated candidates. More sophisticated methods go further by evaluating and selecting promising intermediate generation states during the sampling process (Wu et al., 2023; Li et al., 2024; Kim et al., 2025; Singhal et al., 2025; Ma et al., 2025). While these methods avoid the need for fine-tuning, they often incur significantly higher inference-time costs due to repeated sampling. Although they can boost reward in some cases, they do not directly improve the underlying generative model. In this respect, training-free methods are orthogonal and complementary to fine-tuning-based approaches, and the two can be effectively combined to achieve even stronger performance.
@@ 3 Iterative Distillation Framework for Fine-Tuning Diffusion Models
Our algorithm is motivated by the goal of distilling desirable teacher policies that both maximize task-specific reward functions and preserve the naturalness induced by pre-trained models during fine-tuning. To clarify this motivation, we first define what constitutes teacher policies in our setting, followed by a detailed description of our approach to distilling them.
@@ 3.2 Iterative Distillation
where $u_{t}\in\Delta(\mathcal{X})$ denotes a roll-in distribution, where we will elaborate on in Section 4.1. By simple algebra, up to some constant, this objective reduces to
@@
As we will demonstrate in Section 4.3, this objective can be optimized in practice by approximating value functions and replacing expectations with sample estimates. Notably, the objective is off-policy in nature, meaning that the roll-in distribution $u_{t}$ can be arbitrary and does not need to match the current model policy. Indeed, this objective, commonly referred to as value-weighted maximum likelihood in the RL literature (Peters et al., 2010), has been employed as a scalable and stable approach for offline RL (Peng et al., 2019).
@@
where $\{p^{\mathrm{out}}\}$ denotes the roll-out policy, and $v^{\mathrm{out}}_{t-1}$ is the corresponding soft value function. The roll-out policy is updated in a lazy manner—specifically, it is refreshed every $K$ steps using the current model parameters $p^{\theta_{s}}$. This lazy update scheme is critical for algorithmic stability in the offline setting, preventing rapid changes in the target while still allowing gradual improvement of the student policy toward higher reward. In the next section, we formalize this iterative distillation process as a complete algorithmic procedure.
@@ 4.2 Roll-Out Phase
Recall that soft-value functions in (3) are defined as conditional expectation given $\bar{x}_{t-1}$. While one could estimate these expectations using Monte Carlo sampling or regression, as commonly done in standard RL settings, we recommend a more practical approximation: $\hat{v}_{t-1}(\bar{x}_{t-1}):=r\left(\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})\right)$. Recalling $\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})$ is the denoised prediction from the diffusion model parameterized by the current student policy. This approximation is based on the replacement of the expectation in (3) with its posterior mean. We apply the same approximation to $\hat{v}_{t}(x_{t})$, estimating it via $r(\hat{x}_{0}(x_{t};\theta^{\mathrm{out}}))$.
@@ 4.3 Distillation Phase
Thus far, we have defined the training data distribution via the roll-in phase and specified the supervised signal through the roll-out phase. The final step is to perform distillation by solving a supervised learning problem over the roll-in distribution. Specifically, we minimize the KL divergence between the teacher and student policies, as formalized in (6). In our implementation, soft value functions are approximated by averaging over sampled trajectories. The expectations over roll-in and roll-out distributions are replaced with empirical estimates obtained from samples collected in their respective phases. The resulting loss function is given in (7), which corresponds precisely to a value-weighted maximum likelihood objective.
@@ 5 Comparison with Policy Gradient Methods
In contrast, VIDD optimizes an objective that more closely resembles the forward KL divergence. Since the reverse KL is known to be mode-seeking and can lead to unstable optimization landscapes (Wang et al., 2023; Go et al., 2023; Kim et al., 2025), avoiding reverse KL minimization contributes to more stable and effective fine-tuning. We empirically investigate this effect in the following section.
@@ 6.1.3 Metrics
Recall that our objective is to generate samples with high reward while preserving the naturalness imposed by the pretrained diffusion model in Section 2.2. Accordingly, we report the median reward of generated samples as the primary evaluation metric. As secondary metrics, we include additional measures whose specific choice depends on the task context, such as the naturalness of generated samples. More specifically, we use pLDDT scores for protein design; the 3-mer Pearson correlation (3-mer Corr) between the generated sequences and those in the dataset from Gosai et al. (2023); the negative log-likelihood (NLL) of the generated samples with respect to the pretrained model for small molecule design. For further results, refer to Appendix E.
@@ 6.2 Results
VIDD consistently outperforms baseline methods by achieving higher rewards in $\beta$-sheet content and ipTM binding affinity. Additionally, Figure 3 and Figure 4 illustrate the predicted binder–target complexes and secondary-structure matching tasks, confirming effective binding of the designed binders to their targets and reasonable secondary structures of the designed proteins. Other quantitative results are provided in Appendix E.4, further validating the quality and effectiveness of the generated proteins.
@@
Here, since this reward is technically differentiable, we include an additional baseline, DRAKES (Wang et al., 2024), which directly backpropagates through the reward signal. Notably, as shown in the Pred-Activity column, VIDD outperforms not only DDPP and DDPO, but also DRAKES. In addition to the aforementioned Pred-Activity and 3-mer Corr metrics, we follow Wang et al. (2024) in incorporating ATAC-Acc—an independent binary classification model trained on chromatin accessibility data from the HepG2 cell line (Consortium et al., 2012)—as an orthogonal reward. This is motivated by the fact that Pred-Activity is a trained reward model and thus may be susceptible to overoptimization. In the ATAC-Acc column, our method also exhibits strong performance, suggesting that VIDD would be robust to over-optimization.
@@
VIDD outperforms the fine-tuning baseline methods (DDPO, DDPP, and SFT) in terms of reward. For more quantitative metrics that describe naturalness in molecules, refer to Appendix E.3.
@@ Appendix B Connecting VIDD Objective with KL Divergence
$\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{x_{t}\sim u_{t}}[\mathrm{KL}(p^{\star}_{t-1}(\cdot|x_{t})\|p^{\theta}_{t-1}(\cdot|x_{t}))]$
With simple algebra, this is equal to the following:
$-\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]\right]$
$+\sum_{t}\mathbb{E}_{x_{t}\sim u_{t},x_{t-1}\sim p^{\star}_{t-1}(x_{t-1}|x_{t})}[\log p^{\star}_{t-1}(x_{t-1}|x_{t})].$
Hence, ignoring the second term since this is constant, the objective function reduces to
$\mathop{\mathrm{argmax}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]\right]$
In practice, with slight abust of notation, putting $\exp(v_{t}(x_{t})/\alpha)$ into the roll-in distribution $u_{t}$, we use the following objective:
$\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]\right].$
@@ C.1 Broader Impact
C.2 Limitations
The success of reward-guided fine-tuning such as reinforcement learning critically depends on the quality of the reward signal. However, in reality, reward functions are often imperfect: they may reflect proxy objectives that are only loosely correlated with real-world biological or chemical utility. Poorly designed rewards can lead to over-optimization and exploit the reward without producing truly meaningful or functional outputs. Furthermore, post-training is often sensitive to the precision of the reward—when the reward signal is noisy or misaligned, learning can become unstable or entirely fail. In practice, reward evaluation can also be expensive (e.g., involving structure prediction or simulation), making it essential to design reward mechanisms that are not only accurate but also sample-efficient, enabling effective training under limited reward budget.

### v3

# arXiv:2507.00445v3[cs.LG] 28 Feb 2026

@@ Abstract
We address the problem of fine-tuning diffusion models for reward-guided generation in biomolecular design. While diffusion models have proven highly effective in modeling complex, high-dimensional data distributions, real-world applications often demand more than high-fidelity generation, requiring optimization with respect to potentially non-differentiable reward functions such as physics-based simulation or rewards based on scientific knowledge. Although RL methods have been explored to fine-tune diffusion models for such objectives, they often suffer from instability, low sample efficiency, and mode collapse due to their on-policy nature. In this work, we propose an iterative distillation-based fine-tuning framework that enables diffusion models to optimize for arbitrary reward functions. Our method casts the problem as policy distillation: it collects off-policy data during the roll-in phase, simulates reward-based soft-optimal policies during roll-out, and updates the model by minimizing the KL divergence between the simulated soft-optimal policy and the current model policy. Our off-policy formulation, combined with KL divergence minimization, enhances training stability and sample efficiency compared to existing RL-based methods. Empirical results demonstrate the effectiveness and superior reward optimization of our approach across diverse tasks in protein, small molecule, and regulatory DNA design. The source code is released at (https://divelab.github.io/VIDD/).
@@ 1 Introduction
To address the aforementioned challenges, we propose a new framework, VIDD (Value-guided Iterative Distillation for Diffusion models), designed to maximize possibly non-differentiable rewards in a stable and effective manner. The core idea is to iteratively distill soft-optimal policies—including, which serve as target denoising processes that optimize the reward while remaining close to the current fine-tuned model, as visualized in Figure 1. Concretely, the algorithm proceeds in three iterative steps: (1) roll-in using sufficiently exploratory off-policy trajectories, (2) roll-out to simulate soft-optimal policies, and (3) update the fine-tuned model by minimizing the KL divergence between the soft-optimal and current model policies. Importantly, in (2) and (3), our algorithm effectively leverages value functions tailored to diffusion models to guide fine-tuning, analogous to value-weighted MLE in RL (Peters et al., 2010). Notably, our framework leverages off-policy roll-ins—decoupling data collection from policy updates—and employs forward KL minimization rather than reverse KL, both of which contribute to improved optimization stability over complex reward landscapes.
Our contributions are summarized as follows. We propose a novel algorithm, VIDD, for fine-tuning diffusion models through iterative distillation of target policies composed of both value functions and the current policy. Unlike direct reward backpropagation methods, which require differentiable reward signals, our approach can handle non-differentiable rewards commonly encountered in scientific domains. Algorithmically, our key innovation lies in effectively incorporating value functions specifically tailored to the diffusion models. Empirically, we validate the effectiveness of our method across a range of scientific tasks, particularly in protein and small molecular design.
@@ 1.1 Related Works
When reward functions are differentiable, as is often the case in computer vision, state-of-the-art methods achieve strong performance by directly backpropagating gradients induced by the reward functions on the diffusion model (Clark et al., 2023; Prabhudesai et al., 2023; Wu et al., 2024; Wang et al., 2024). However, in many scientific domains such as biology and molecule field, reward functions are often non-differentiable (Lisanza et al., 2024; Hie et al., 2022). In such settings, a natural approach is to adopt RL techniques that optimize reward without requiring differentiability (Zhang et al., 2024; Gupta et al., 2025; Ektefaie et al., 2024; Vogt et al., 2024; Deng et al., 2024; Venkatraman et al., 2024; Rector-Brooks et al., 2024; Zekri and Boullé, 2025). For instance, DPOK (Fan et al., 2023) and DDPO (Black et al., 2024) adapt PPO, a stabilized variant of policy gradient methods, to the diffusion model fine-tuning. However, these methods often suffer from instability when applied to diffusion models (Clark et al., 2023, Figure 3), due to their inherently on-policy nature and the use of a reverse KL divergence objective between the fine-tuned policy and a target policy (discussed further in Section 5). In contrast, our method avoided on-policy updates and instead leverages a forward KL divergence objective, which stabilizes training and mitigates mode collapse.
@@
An alternative line of work focuses on training-free techniques that aim to improve rewards solely at inference time. The most straightforward example is Best-of-N sampling, which selects the highest-reward output from N generated candidates. More sophisticated methods go further by evaluating and selecting promising intermediate generation states during the sampling process (Wu et al., 2023; Li et al., 2024; Kim et al., 2025; Singhal et al., 2025; Ma et al., 2025; Tang et al., 2025). While these methods avoid the need for fine-tuning, they often incur significantly higher inference-time costs due to repeated sampling. Although they can boost reward in some cases, they do not directly improve the underlying generative model. In this respect, training-free methods are orthogonal and complementary to fine-tuning-based approaches, and the two can be effectively combined to achieve even stronger performance.
@@ 3 Iterative Distillation Framework for Fine-Tuning Diffusion Models
Our algorithm is motivated by the goal of distilling desirable teacher policies that maximize task-specific reward functions. To clarify this motivation, we first define what constitutes teacher policies in our setting, followed by a detailed description of our approach to distilling them.
@@ 3.2 Iterative Distillation
where $u_{t}\in\Delta(\mathcal{X})$ denotes a roll-in distribution, where we will elaborate on in Section 4.1. By simple algebra (see the detailed derivation in Appendix B), up to some constant, this objective reduces to
@@
As we will demonstrate in Section 4.3, this objective can be optimized in practice by approximating value functions and replacing expectations with sample estimates. Notably, the objective is off-policy in nature, meaning that the roll-in distribution $u_{t}$ can be arbitrary and does not need to match the current model policy. Indeed, this objective, commonly referred to as value-weighted maximum likelihood in the RL literature (Peters et al., 2010), has been employed as a scalable and stable approach for off-policy RL (Peng et al., 2019).
@@
where $\{p^{\mathrm{out}}\}$ denotes the roll-out policy, and $v^{\mathrm{out}}_{t-1}$ is the corresponding soft value function. The roll-out policy is updated in a lazy manner—specifically, it is refreshed every $K$ steps using the current model parameters $p^{\theta_{s}}$. This lazy update scheme is critical for algorithmic stability in the off-policy setting, preventing rapid changes in the target while still allowing gradual improvement of the student policy toward higher reward. In the next section, we formalize this iterative distillation process as a complete algorithmic procedure.
@@ 4.2 Roll-Out Phase
Recall that soft-value functions in (3) are defined as conditional expectation given $\bar{x}_{t-1}$. While one could estimate these expectations using Monte Carlo sampling or regression, as commonly done in standard RL settings, we recommend a more practical approximation: $\hat{v}_{t-1}(\bar{x}_{t-1}):=r\left(\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})\right)$. Recalling $\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})$ is the denoised prediction from the diffusion model parameterized by the current student policy. This approximation is based on the replacement of the expectation in (3) with its posterior mean. We apply the same approximation to $\hat{v}_{t}(x_{t})$, estimating it via $r(\hat{x}_{0}(x_{t};\theta^{\mathrm{out}}))$. Further discussions on soft value functions can be found in Appendix F.
@@ 4.3 Distillation Phase
Thus far, we have defined the training data distribution via the roll-in phase and specified the supervised signal through the roll-out phase. The final step is to perform distillation by solving a supervised learning problem over the roll-in distribution. Specifically, we minimize the KL divergence between the teacher and student policies, as formalized in (6). In our implementation, soft value functions are approximated using estimates from trajectories. The expectations over roll-in and roll-out distributions are replaced with empirical estimates obtained from samples collected in their respective phases. The resulting loss function is given in (7), which corresponds precisely to a value-weighted maximum likelihood objective.
@@ 5 Comparison with Policy Gradient Methods
In contrast, VIDD optimizes an objective that more closely resembles the forward KL divergence. Since the reverse KL is known to be mode-seeking and can lead to unstable optimization landscapes (Wang et al., 2023; Go et al., 2023; Kim et al., 2025), avoiding reverse KL minimization contributes to more stable and effective fine-tuning.
@@ 6.1.3 Metrics
Recall that our objective is to generate samples with high desired reward as in Section 2.2. Accordingly, we report the median reward of generated samples as the primary evaluation metric. As secondary metrics, we include additional measures whose specific choice depends on the task context, such as the naturalness of generated samples. More specifically, we use pLDDT scores for protein design; the 3-mer Pearson correlation (3-mer Corr) between the generated sequences and those in the dataset from Gosai et al. (2023); the negative log-likelihood (NLL) of the generated samples with respect to the pretrained model for small molecule design. For further results, refer to Appendix E.
@@ 6.2 Results
VIDD consistently outperforms baseline methods by achieving higher rewards in $\beta$-sheet content and ipTM binding affinity. Additionally, Figure 3 and Figure 4 illustrate the predicted binder–target complexes and secondary-structure matching tasks, confirming effective binding of the designed binders to their targets and reasonable secondary structures of the designed proteins. Other quantitative results are provided in Appendix E.4, further validating the quality and effectiveness of the generated proteins, as well as effects of varying parameters such as mixture of roll-in policy, lazy update interval and regularization coefficient.
@@
Here, since this reward is technically differentiable, we include an additional baseline, DRAKES (Wang et al., 2024), which directly backpropagates through the reward signal. Notably, as shown in the Pred-Activity column, VIDD outperforms not only DDPP and DDPO, but also DRAKES. In addition to the aforementioned Pred-Activity and 3-mer Corr metrics, we follow Wang et al. (2024) in incorporating ATAC-Acc—an independent binary classification model trained on chromatin accessibility data from the HepG2 cell line (Consortium and others, 2012)—as an orthogonal reward. This is motivated by the fact that Pred-Activity is a trained reward model and thus may be susceptible to overoptimization. In the ATAC-Acc column, our method also exhibits strong performance, suggesting that VIDD would be robust to over-optimization. More quantitative results can be found in Appendix E.2.
@@
VIDD outperforms the fine-tuning baseline methods in terms of reward. More quantitative metrics that describe naturalness in molecules is in Appendix E.3.
@@ Appendix B Connecting VIDD Objective with KL Divergence
$\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{x_{t}\sim u_{t}}[\mathrm{KL}(p^{\star}_{t-1}(\cdot|x_{t})\|p^{\theta}_{t-1}(\cdot|x_{t}))].$
Based on the definition of conditional KL divergence
$\mathrm{KL}\bigl(p(\cdot\mid x_{t})\,\|\,q(\cdot\mid x_{t})\bigr)$ $=\mathbb{E}_{x_{t-1}\sim p(\cdot\mid x_{t})}\bigl[\log p(x_{t-1}\mid x_{t})-\log q(x_{t-1}\mid x_{t})\bigr],$
and the definition of soft optimal policy $p^{\star}_{t-1}$:
$p^{\star}_{t-1}(\cdot|x_{t})$ $=\frac{p^{\mathrm{pre}}_{t-1}(\cdot|x_{t})\exp(v_{t-1}(\cdot)/\alpha)}{\exp(v_{t}(x_{t})/\alpha)},$
our objective can be rewritten as
$\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{x_{t}\sim u_{t},x_{t-1}\sim p^{\star}_{t-1}(x_{t-1}|x_{t})}\Bigl[\log p^{\star}_{t-1}(x_{t-1}\mid x_{t})-\log p^{\theta}_{t-1}(x_{t-1}\mid x_{t})\Bigr]$
$=$ $\mathop{\mathrm{argmin}}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t},x_{t-1}\sim p^{\star}_{t-1}(x_{t-1}|x_{t})}[\log p^{\star}_{t-1}(x_{t-1}|x_{t})]$
$-\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]\right].$
Hence, ignoring the first term since this is constant, the objective function reduces to
$\mathop{\mathrm{argmax}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]\right],$
thus we obtain (5).
@@ C.1 Broader Impact
C.2 Limitations and future works
The success of reward-guided fine-tuning such as reinforcement learning critically depends on the quality of the reward signal. However, in reality, reward functions are often imperfect: they may reflect proxy objectives that are only loosely correlated with real-world biological or chemical utility. Poorly designed rewards can lead to over-optimization and exploit the reward without producing truly meaningful or functional outputs.
Furthermore, post-training is often sensitive to the precision of the reward—when the reward signal is noisy or misaligned, learning can become unstable or entirely fail. In practice, reward evaluation can also be expensive (e.g., involving structure prediction or simulation), making it essential to design reward mechanisms that are not only accurate but also sample-efficient, enabling effective training under limited reward budget.
Finally, since reward signals are imperfect and often expensive to evaluate, hyperparameters that regulate reward exploitation such as rollout frequency, may require some tuning in practice. Overly aggressive configurations can amplify reward noise, while overly conservative choices may cap achievable gains, reflecting a practical tradeoff inherent to reward-driven optimization. Designing fully adaptive schedules for these parameters would require domain-specific assumptions about reward reliability and exploration metrics, which falls outside the scope of this work but represents a promising engineering extension built on top of our framework.